OpenStudy (anonymous):

Evaluate the integral: 5x^3 ln

OpenStudy (zzr0ck3r):

? 5x^3*ln(x)?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

do it by parts

OpenStudy (zzr0ck3r):

let u = ln(x) and dv = x^3dx

OpenStudy (anonymous):

i have u=lnx and dv=5x^3 but i dont know how to solve and yes i know the formula for it am just having trouble

OpenStudy (zzr0ck3r):

whats the formula?

OpenStudy (anonymous):

uv-(swirl)vdu

OpenStudy (zzr0ck3r):

x du = dx and v = (5/4)x^4

OpenStudy (anonymous):

x du = dx?

OpenStudy (zzr0ck3r):

ln(x)*(5/4)x^4 - int(((5/4)*x^4)/x)dx

OpenStudy (zzr0ck3r):

du = (1/x) dx

OpenStudy (zzr0ck3r):

so ln(x)(5/4)*x^4 - (5/4)*int(x^3 dx)

OpenStudy (zzr0ck3r):

so ln(x) (5/4)x^4- (5/4)(1/4)x^4 + c

OpenStudy (zzr0ck3r):

5*x^4*ln(x)/4-((5*x^4)/16) + c

OpenStudy (anonymous):

would be the final answer?

OpenStudy (zzr0ck3r):

sure it simplifies to (5/16)x^4(4ln(x)-1)

OpenStudy (zzr0ck3r):

does it make sense?

OpenStudy (anonymous):

yess awesome thank you for your help! could you also please help me with another if you have time?

OpenStudy (zzr0ck3r):

sure post it on another thread just in case i cant

OpenStudy (anonymous):

okay hang on